OpenStudy (anonymous):

What is the solution to the rational equation x over x squared minus 9 plus 1 over x plus 3 equals 1 over 4x minus 12 ? x = -9 over 7 x = 15 x = 15 over 7 x = 9 i will reward a medal

OpenStudy (anonymous):

\[\frac{ x }{ x ^{2}-9 }+\frac{ 1 }{ x+3 }=\frac{ 1 }{ 4x-12 }\]

OpenStudy (anonymous):

that equation ubove is the equation for the question